OpenStudy (pokemon23):

Mildred brought 1 cent stamps,32 cent stamps, and 33 cent stamps for $21.80. The number of 1 cent stamps exceed the number of 33 cent stamps by 50. The number of 32 cents was 10 less than twice the number of 33 cents stamps. How many of each kind did she buy? Mildred brought 1 cent stamps,32 cent stamps, and 33 cent stamps for $21.80. The number of 1 cent stamps exceed the number of 33 cent stamps by 50. The number of 32 cents was 10 less than twice the number of 33 cents stamps. How many of each kind did she buy? @Mathematics

OpenStudy (pokemon23):

HELP ME GET AN EQUATION

OpenStudy (pokemon23):

thanks ashe

OpenStudy (pokemon23):

and mert

OpenStudy (mertsj):

x = Number of 33 cent stamps x+50 = Number of 1 cent stamps 2x-10 = number of 32 cent stamps 33x = value of 33 cent stamps 1(x+50) = value of 1 cent stamps 32(2x-10) = value of 32 cent stamps 33x+(x+50) + 32(2x-10) = 2180 All values in CENTS

OpenStudy (ash2326):

ok let no. of 1 cent stamps be x, no. of 32 cent stamps be y and z be the 33 cent stamps 1st eq. x*1+y*32+z*33=2180; 2nd eq. x-z=50; and third is 2z-y=10; u can solve these 3 equations to get x , y and z

OpenStudy (mertsj):

33x + x +50 + 64x -320 = 2180 98x - 270 = 2180 98x = 2450 x = 25= number of 33 cent stamps x+50 = 75 = number of 1 cent stamps 2x-10 = 40 = number of 32 cent stamps